A capacitor consists of two closely spaced metal conductors of large area, separated by a thin insulating foil. It has an electrical capacity of 3600.0 μF and is charged to a potential difference of 78.0 V. Calculate the amount of energy stored in the capacitor.

_____???

Calculate the charge on this capacitor when the electrical energy stored in the capacitor is 10.95 J.

_____???

If the two plates of the capacitor have their separation increased by a factor of 6 while the charge on the plates remains constant, by what factor is the energy stored in the capacitor increased?

______???

Please help me solve these! Thanks!
Answer #1

1) = (1/2) × (3600×10-6)× 782

= 10.95 J

2)

Q = (3600×10-6) × 78 = 0.281 C

3) as

Energy is given by

As d becomes 6 times , Therefore energy increases by factor of 6